OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sine of seven pi divided by eight.

OpenStudy (anonymous):

\[\frac{7\pi}{8}\] is half of \[\frac{7\pi}{6}\]

OpenStudy (anonymous):

thats it?

Miracrown (miracrown):

@satellite73 7π/8 is NOT half of 7π/6

Miracrown (miracrown):

@jon321 do you know what the half of 7π/8 would be then?

OpenStudy (anonymous):

1.37

OpenStudy (anonymous):

lol so it isnt!

Miracrown (miracrown):

nope lol

OpenStudy (anonymous):

multiply \(\frac{7\pi}{8}\) by \(2\)

OpenStudy (anonymous):

and forget about decimals, you are supposed to use the "half angle" formula, like it says

OpenStudy (anonymous):

.3827

OpenStudy (anonymous):

7pie/4

OpenStudy (anonymous):

yeah that one

OpenStudy (anonymous):

thats the answer?

jimthompson5910 (jim_thompson5910):

you'll need to know the value of cos(7pi/4) and you'll plug it into the half angle identity see this reference sheet http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf specifically, the "Half Angle Formulas" section at the top of page 2